Circular No. 6304

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 B2 (HYAKUTAKE) Total visual magnitude estimates: Feb. 1.49 UT, 10.1: (A. Hale, Cloud-

croft, NM, 0.41-m reflector; coma diameter ∼ 2′.5); 1.73, 8.9 (D. A. J. Seargent, The Entrance, N.S.W., 25 × 100 binoculars; coma 4′); 1.81, 11.7 (Y. Kushida, Yatsugatake, Japan, 0.40-m reflector; coma 2′.5–3′.0).

The orbital elements on IAUC 6303 indicate that the comet will pass only 0.11 AU from the earth on Mar. 26 on its way to perihelion rather more than a month later. The comet can be expected to become a nakedeye object, both on account of the approach to the earth, and later as it approaches perihelion. The predicted magnitude in the following ephemeris for around those times is, of course, highly uncertain. There is also a very large positional uncertainty, particularly in late March, when it amounts to tens of degrees.

1996 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Jan. 28 14^{h}29.^{m}01 −24^{o}53.^{′}4 1.953 2.102 84.^{o}6 27.^{o}8 10.2
Feb. 2 14 33.06 −24 56.3 1.788 2.022  88.7  29.2  9.8
7 14 36.94 −24 54.9 1.623 1.941  92.9  30.5  9.4
12 14 40.63 −24 47.7 1.456 1.858  97.1  31.8  9.0
17 14 44.07 −24 32.9 1.289 1.773 101.4  33.1  8.5
22 14 47.21 −24 07.0 1.122 1.687 105.9  34.3  8.0
27 14 49.98 −23 24.8 0.955 1.599 110.5  35.5  7.4
Mar. 3 14 52.28 −22 16.6 0.789 1.508 115.4  36.4  6.8
8 14 53.95 −20 23.1 0.623 1.416 120.6  37.1  6.0
13 14 54.65 −16 58.2 0.460 1.321 126.5  37.2  5.0
18 14 53.46 − 9 42.4 0.302 1.223 133.5  36.2  3.8
23 14 46.46 +11 55.0 0.159 1.122 139.0  35.6  2.0
28 12 28.3 +80 59.2 0.118 1.017  96.0  77.4  0.9
Apr. 2 3 24.64 +57 28.7 0.236 0.909  60.9 106.0  1.9
7 3 12.26 +46 13.5 0.387 0.796  48.0 110.8  2.4
12 3 06.26 +40 56.1 0.545 0.677  39.5 109.7  2.5
17 3 00.45 +37 22.2 0.706 0.554  32.2 105.0  2.2
22 2 53.20 +34 04.2 0.869 0.426  25.0  95.7  1.5
27 2 43.52 +29 55.5 1.033 0.303  17.0  76.6  0.4

COMET 73P/SCHWASSMANN-WACHMANN 3 Corrigendum. On IAUC 6301, lines 10 and 26, for Häufl read Käufl.

1996 February 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Brian G. Marsden